# On the Supersingular Locus of the $\operatorname{GU}(2,n-2)$ Shimura Variety

### v1

@@ 1 Introduction
Let $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$. Then $\pi(X_{w_{k,l}}(b))$ is a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $\pi(X_{w^{\prime}_{k,l}}(b))$. In particular, if $k+l\geq n+3$ (resp. $k+l\leq n+2$), then $\pi(X_{w_{k,l}}(b))$ is an iterated fibration of rank $\lfloor\frac{k+l-n-1}{2}\rfloor+\lfloor\frac{k-1}{2}\rfloor$ (resp. $\frac{k-1}{2}$) over $\pi(X_{w_{1,n-k+2}}(b))$ (resp. $\pi(X_{w_{1,l}}(b))$), whose fibers are all $\mathbb{A}^{1}$.
@@ 2.4 Length Positive Elements
Thanks to Kottwitz [25], a $\sigma$-conjugacy class $[b]\in B(G)$ is uniquely determined by two invariants: the Kottwitz point $\kappa(b)\in\pi_{1}(G)/((1-\sigma)\pi_{1}(G))$ and the Newton point $\nu_{b}\in X_{*}(T)_{\mathbb{Q},+}$. Clearly, $X_{w}(b)=\emptyset$ if $\kappa(b)\neq\kappa(\dot{w})$. We say that $b\in G(L)$ is basic if $\nu_{b}$ is central. The following theorem is a refinement of the non-emptiness criterion in [13], which is conjectured by Lim in [28] and proved by Schremmer in [31, Proposition 5].
@@ 3.2 Computation in the Affine Weyl Group
If this is the case, $X_{w_{k,l}}(\tau)$ is $\mathbb{J}_{\tau}$-equivariant universally homeomorphic to a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $X_{w^{\prime}_{k,l}}(\tau)$. In particular, if $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$ and $k+l\leq n+2$ (resp. $k+l\geq n+3$), then $X_{w_{k,l}}(\tau)$ is an iterated fibration of rank $\lfloor\frac{k+l-n-1}{2}\rfloor+\lfloor\frac{k-1}{2}\rfloor$ (resp. $\frac{k-1}{2}$) over $X_{w_{1,l}}(\tau)$ (resp. $X_{w_{1,n-k+2}}(\tau)$).
@@ 4.1 Iterated Fibrations over Deligne-Lusztig Varieties
Let $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$. Then $\pi(X_{w_{k,l}}(b))$ is $G(F)$-equivariant universally homeomorphic to a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $\pi(X_{w^{\prime}_{k,l}}(b))$. In particular, if $k+l\geq n+3$ (resp. $k+l\leq n+2$), then
@@
and $\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{k,l}))$ is an irreducible component, which is $G(\mathcal{O}_{F})$-equivariant universally homeomorphic to an iterated fibration of rank $\lfloor\frac{k+l-n-1}{2}\rfloor+\lfloor\frac{k-1}{2}\rfloor$ (resp. $\frac{k-1}{2}$) over $\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,n-k+2}))$ (resp. $\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,l}))$).
@@
Let $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$. By Lemma 4.2 and Proposition 2.6, $\pi_{S(w_{k,l},\sigma)}(X_{w_{k,l}}(b))$ is $G(F)$-equivariant universally homeomorphic to a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $\pi_{S(w_{k,l},\sigma)}(X_{w^{\prime}_{k,l}}(b))$. In particular, if $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$ and $k+l\geq n+3$ (resp. $k+l\leq n+2$), then $\pi_{S(w_{k,l},\sigma)}(X_{w_{k,l}}(b))$ is an iterated fibration of rank $\lfloor\frac{k+l-n-1}{2}\rfloor+\lfloor\frac{k-1}{2}\rfloor$ (resp. $\frac{k-1}{2}$) over $\pi_{S(w_{k,l},\sigma)}(X_{w_{1,n-k+2}}(b))$ (resp. $\pi_{S(w_{k,l},\sigma)}(X_{w_{1,l}}(b))$). By Lemma 4.1, we have $\operatorname{supp}_{\sigma}(w_{k,l})_{1}\cup S(w_{k,l},\sigma)_{1}=S$ if $k=1$ and $l\geq\frac{n+3}{2}$. Thus the last statement follows from these results and the isomorphism $\pi_{S(w_{k,l},\sigma)}(X_{w_{k,l}}(b))\cong\pi(X_{w_{k,l}}(b))$. This finishes the proof. ∎
@@
If $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\operatorname{DL}}$, then Theorem C holds by setting $P_{k,l}=P_{\operatorname{supp}_{\sigma}(w_{k,l})_{1}\cup S(w_{k,l},\sigma)_{1}}$ and $X_{k,l}=Y_{k,l}=\tau_{1}\pi(Y(w_{k,l}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{k,l}))$. If $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$ and $k+l\geq n+3$ (resp. $k+l\leq n+2$), then Theorem C holds by setting $P_{k,l}=G(\mathcal{O})$, $X_{k,l}=\tau_{1}\pi(Y(w_{1,n-k+2}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,n-k+2}))$ (resp. $X_{k,l}=\tau_{1}\pi(Y(w_{1,l}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,l}))$) and $Y_{k,l}=\tau_{1}\pi(Y(w_{k,l}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{k,l}))$.

### v2

@@ 1 Introduction
Let $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$. Then $\pi(X_{w_{k,l}}(b))$ is a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $\pi(X_{w^{\prime}_{k,l}}(b))$. In particular, if $k+l\leq n+2$ (resp. $k+l\geq n+3$), then $\pi(X_{w_{k,l}}(b))$ is an iterated fibration of rank $\frac{k-1}{2}$ (resp. $k+\frac{l-n-3}{2}$) over $\pi(X_{w_{1,l}}(b))$ (resp. $\pi(X_{w_{1,n-k+2}}(b))$), whose fibers are all $\mathbb{A}^{1}$.
@@ 2.4 Length Positive Elements
Thanks to Kottwitz [25], a $\sigma$-conjugacy class $[b]$ of $b\in G(L)$ is uniquely determined by two invariants: the Kottwitz point $\kappa(b)\in\pi_{1}(G)/((1-\sigma)\pi_{1}(G))$ and the Newton point $\nu_{b}\in X_{*}(T)_{\mathbb{Q},+}$. Clearly, $X_{w}(b)=\emptyset$ if $\kappa(b)\neq\kappa(\dot{w})$. We say that $b\in G(L)$ is basic if $\nu_{b}$ is central. The following theorem is a refinement of the non-emptiness criterion in [13], which is conjectured by Lim in [28] and proved by Schremmer in [31, Proposition 5].
@@ 3.2 Computation in the Affine Weyl Group
If this is the case, $X_{w_{k,l}}(\tau)$ is $\mathbb{J}_{\tau}$-equivariant universally homeomorphic to a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $X_{w^{\prime}_{k,l}}(\tau)$. In particular, if $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$ satisfies (i) (resp. (ii)), then $X_{w_{k,l}}(\tau)$ is an iterated fibration of rank $\frac{k-1}{2}$ (resp. $k+\frac{l-n-3}{2}$) over $X_{w_{1,l}}(\tau)$ (resp. $X_{w_{1,n-k+2}}(\tau)$).
@@ 4.1 Iterated Fibrations over Deligne-Lusztig Varieties
Let $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$. Then $\pi(X_{w_{k,l}}(b))$ is $G(F)$-equivariant universally homeomorphic to a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $\pi(X_{w^{\prime}_{k,l}}(b))$. In particular, if $k+l\leq n+2$ (resp. $k+l\geq n+3$), then
@@
and $\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{k,l}))$ is an irreducible component, which is $G(\mathcal{O}_{F})$-equivariant universally homeomorphic to an iterated fibration of rank $\frac{k-1}{2}$ (resp. $k+\frac{l-n-3}{2}$) over $\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,l}))$ (resp. $\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,n-k+2}))$).
@@
Let $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$. By Lemma 4.2 and Proposition 2.6, $\pi_{S(w_{k,l},\sigma)}(X_{w_{k,l}}(b))$ is $G(F)$-equivariant universally homeomorphic to a Zariski-locally trivial $\mathbb{A}^{1}$-bundle over $\pi_{S(w_{k,l},\sigma)}(X_{w^{\prime}_{k,l}}(b))$. In particular, if $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$ satisfies (i) (resp. (ii)) of Theorem 3.4, then $\pi_{S(w_{k,l},\sigma)}(X_{w_{k,l}}(b))$ is an iterated fibration of rank $\frac{k-1}{2}$ (resp. $k+\frac{l-n-3}{2}$) over $\pi_{S(w_{k,l},\sigma)}(X_{w_{1,l}}(b))$ (resp. $\pi_{S(w_{k,l},\sigma)}(X_{w_{1,n-k+2}}(b))$). By Lemma 4.1, we have $\operatorname{supp}_{\sigma}(w_{k,l})_{1}\cup S(w_{k,l},\sigma)_{1}=S$ if $k=1$ and $l\geq\frac{n+3}{2}$. Thus the last statement follows from these results and the isomorphism $\pi_{S(w_{k,l},\sigma)}(X_{w_{k,l}}(b))\cong\pi(X_{w_{k,l}}(b))$. This finishes the proof. ∎
@@
If $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\operatorname{DL}}$, then we set $P_{k,l}=P_{\operatorname{supp}_{\sigma}(w_{k,l})_{1}\cup S(w_{k,l},\sigma)_{1}}$ and $X_{k,l}=Y_{k,l}=\tau_{1}\pi(Y(w_{k,l}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{k,l}))$. If $w_{k,l}\in{{}^{S}\mathrm{Adm}}(\mu)_{\neq\operatorname{DL}}$ and $k+l\leq n+2$ (resp. $k+l\geq n+3$), then we set $P_{k,l}=G(\mathcal{O})$, $X_{k,l}=\tau_{1}\pi(Y(w_{1,l}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,l}))$ (resp. $X_{k,l}=\tau_{1}\pi(Y(w_{1,n-k+2}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{1,n-k+2}))$) and $Y_{k,l}=\tau_{1}\pi(Y(w_{k,l}))\cong\tau_{1}\pi_{S(w_{k,l},\sigma)}(Y(w_{k,l}))$. This notation justifies Theorem C.